# Revised bounds on local cosmic strings from NANOGrav observations

### v1

@@ Abstract
In a recent paper, the NANOGrav collaboration studied new physics explanations of the observed pulsar timing residuals consistent with a stochastic gravitational wave background (SGWB) [1], including cosmic strings in the Nambu-Goto (NG) approximation. Analysing one of current models for the loop distribution, it was found that the cosmic string model is disfavored compared to other sources, for example, super massive black hole binaries (SMBHBs). When both SMBHB and cosmic string models are included in the analysis, an upper bound on a string tension $G\mu\lesssim 10^{-10}$ was derived. However, the analysis did not accommodate results from cosmic string simulations in an underlying field theory, which indicate that at most a small fraction of string loops survive long enough to emit GW. Following and extending our previous study [2], we suppose that a fraction $f_{\rm NG}$ of string loops follow NG dynamics and emit only GWs, and study the three different models of the loop distribution discussed in the LIGO-Virgo-KAGRA (LVK) collaboration analyses. We re-analyze the NANOGrav 15yrs data with our signal models by using the NANOGrav ENTERPRISE analysis code via the wrapper PTArcade. We find that loop distributions similar to LVK Model B and C yield higher Bayes factor than Model A analyzed in the NANOGrav paper, as they can more easily accommodate a blue-tilted spectrum of the observed amplitude. Furthermore, because of the degeneracy of $G\mu$ and $f_{\rm NG}$ in determining the signal amplitude, our posterior distribution extends to higher values of $G\mu$, and in some cases the uppermost value of credible intervals is close to the Cosmic Microwave Background limit $G\mu\lesssim 10^{-7}$. Hence, in addition to the pulsar timing array data, further information about the fraction of long-lived loops in a cosmic string network is required to constrain the string tension.
@@ 1 Introduction
Recently, PTA collaborations (NANOGrav, EPTA/InPTA, PPTA and CPTA) have reported evidence for an isotropic SWGB as the expected Hellings-Downs (HD) angular correlation between pulsars’ line of sight were measured at $3-4\sigma$ confidence level [20, 21, 22, 23]. With high significance of the common spectrum and the strong evidence of HD correlation, it is highly motivated to investigate whether current observations can be explained by SGWB from the population of cosmic string loops [1, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] (see also Refs. [40, 41] for global strings.). Indeed, NANOGrav and EPTA performed Bayesian inference for the models of NG loop distribution [42, 43], whose SGWB spectrum is controlled by the dimensionless string tension $G\mu$. Because the NG models have difficulty reproducing the blue-tilted spectrum favored by the data with consistent amplitude, upper bounds on the string tension are derived when combining the NG models and the “more favored” super massive black hole binary (SMBHB) signal [1, 24].
@@ 2.1 SGWB from field theory string loops
As mentioned in the introduction, long-lived oscillating NG-like loops, required for a significant GW signal, have not yet been observed in large-scale field theory simulations [52]. Therefore their length distribution, needed for calculations of the GW power spectrum, is completely unknown. In Ref. [52], (a part of) this uncertainty in their distribution was parameterised by allowing a fraction $f_{\rm NG}$ of loops to survive to radiate only gravitationally. It was assumed that the NG-like loops would have the same length distribution as in an NG network ${\sf n}(l,t)$, and hence that the distribution of NG-like loops would be $f_{\rm NG}{\sf n}(l,t)$. Then the SGWB from NG-like distribution in the AH string network is quantified as
@@
In the rest of the section, we discuss possible models of NG loop distributions.
We emphasise that use of the models of NG loop distribution is only a first approximation. Since there is no information from the AH simulations what ${\sf n}(l,t)$ should be used, the interpretation of $f_{\rm NG}$ as parameterising the fraction of NG-like loops depends on how close the NG models are to the true distribution of any long-lived loops in a field theory network. The development of improved models of loop distribution is left to future work, with the expectation that the correction in Eq. (2.1) is O(1) [^1] For example, assuming that the loop size at the production well agrees with the BOS model, the authors of Ref. [64] quantify the suppression of the efficiency of loop production and the resultant overall amplitude of SGWB due to the classical radiation, which was only by a factor $\sim$ 2.. In any case, the difference in the following results for different reference NG loop models should be understood as the underlying uncertainty in the true loop distribution
@@ 2.2 SGWB from NG string loops
with $H(z)$ and $t(z)$ being the Hubble parameter and cosmic time at redshift $z$. The function $C_{n}(f)$ gives the number density of loops emitting GWs observed at frequency $f$ in the $n$-th harmonic, while the average loop power spectrum $P_{n}$ represents the average GW power emitted by the $n$-th harmonic of a loop. The constants $P_{n}$ depend on the average shape of the loops, and in particular the number of cusps per oscillation and the number of kinks travelling around the loop, which dictate the high frequency behaviour. Their sum (or the so-called decay constant of strings) $\Gamma$ can be decomposed into three contributions as
@@ 2.2.1 BOS model (LVK-Model A)
As a power spectrum of loops $P_{n}$, the “smoothed” model [65] is often adopted for this BOS distribution. It is constructed from the numerical simulation taking into account the gravitational backreaction and different from the simple decomposition in Eq. (2.5). Following the results of their simulation, we set $\Gamma=50$ and extract the values of $P_{n}$ from the Figure 3 and 4 of Ref. [65]. As another example in this class of power spectra, we also consider the kink-dominated case $(N_{c},N_{k})=(1,100)$ with Eq. (2.5), which was analyzed in Refs. [19, 66, 21] given the uncertainty on the initial number of kinks at the loop production.
@@ 2.2.2 LRS model (LVK-Model B)
In the context of recent PTA observation, this model was analyzed (together with the BOS model) by the EPTA collaboration with their second data release [24]. Although they do not perform a model selection analysis, the posterior distribution they obtained somewhat supports the NANOGrav result mentioned above. That is, when they include the SMBHB signal in their analysis, they observe the posterior distribution for $G\mu$ extends downward to the limit of the prior (see the right panel of Fig. 8 in App. A or Fig. 15 of Ref. [24]). This indicates that with the addition of the SMBHB signal, which is more preferred by the data, the magnitude of string signal (or $G\mu$) could only be bounded above.
Following the analyses performed in Refs. [66, 24], here we consider two different scenarios for the loop power spectrum. The first one is $(N_{c},N_{k})=(2,0)$, which yields $\Gamma=57$, a value close to that observed in the simulations. Another one is again $(N_{c},N_{k})=(1,100)$ allowing for the large number of kinks.
@@ 3.1 Summary of the analyses
As discussed in Refs. [1, 24], however, the information carried by the “signal” is not particularly constraining at the moment and the result should crucially depend on the priors. Therefore, especially under the significant uncertainty in the underlying AH loop distribution, too much emphasis should not be placed on the exact number of Bayes factors reported in this study. Nevertheless, it can be considered as a good indicator for comparing which of the models we analyzed is relatively better at explaining the data.
@@ 3.2.1 $f_{\rm NG}$-BOS models
We also analyze the case where the indicated SGWB signal is the superposition of the signal from SMBHB and the AH strings. In Fig. 3, the posterior distribution of the string parameters and SMBHB is plotted. Hereafter in the posterior plots, gw-bhb-0 and gw-bhb-1 represent $A_{\rm BHB}$ and $\gamma_{\rm BHB}$ respectively. Again, the left panel is for the smoothed loop power spectrum and the right panel is for the kink dominated spectrum. In both cases, it can be seen that the data favours the sub-dominance of the string signal and the credible intervals of $(G\mu,f_{\rm NG})$ are bounded from above. As expected, the upper most value of $G\mu$ in the credible interval becomes larger than the case with pure BOS model $[-11.99,-9.90]$ (see Table 4 in Ref. [1]). In other words, higher $G\mu$ (or symmetry breaking scales) are to some extent still allowed for by taking into account that not all string loops emit GWs.
@@ 4 Discussion
In quantifying the SGWB signal from long-lived AH string loops, here we took three established models for NG loop distribution [42, 43, 68]. These models are only approximations since there are no long-lived loops observed in simulations of networks the AH model. Therefore, the physical interpretation of $f_{\rm NG}$ as a fraction of long-lived loops depends on how close the model distribution is to the actual loop distribution. In this sense, further study of possible NG-like loop production is indispensable for the refinement of SGWB constraint on the AH model: the current status is only that $f_{\rm NG}\lesssim 0.1$.
@@
Our analyses were based on the wrapper PTArcade [60, 61] and conducted in the similar way to the NANOGrav study of new physics [1]. Note that while the BOS model was analyzed as a NG loop distribution model and subjected to the model selection analysis by NANOGrav [1], the LRS and LVK-Model-C were not (see Ref. [24] for the posterior for the LRS model derived from EPTA data). As summarized in Sec. 3, we performed Bayesian inference and model selection analysis for all three models assuming only the signal from strings and also assuming the signal from SMBHB. We also investigated the different loop power spectra for each model. In all cases our posterior for the AH model extends to a higher value of $G\mu$ (with a decrease in the value of $f_{\rm NG}$), compared to the pure NG model results in Refs. [1, 24]. Phenomenologically, this suggests that symmetry breaking in the AH model is still possible on relatively high energy scales. It is also worth mentioning that for all the loop distribution models, the posterior of SMBHB parameters become more consistent with the theory-motivated prior used in the NANOGrav paper [20], when the string signal is superimposed to the SMBHB signal. When comparing each model, the LRS-like and model-C-like distribution overall had larger Bayes factors (against the SMBHB signals) than the BOS-like distribution. This is due to the fact that for a wider range of $G\mu$ values, loops of the LRS model and LVK model-C produce a blue-tilted SGWB, which is favored by the data. We should, however, note that the differences in our results for different reference NG loop models should be regarded as an estimate of the uncertainty in the AH loop distribution.
@@
After deriving the posterior distributions, we compared them to the existing bounds such as $G\mu\lesssim 10^{-7}$ from CMB [75, 55] and $f_{\rm NG}\lesssim 0.1$ derived from the AH loop simulation [52]. For the BOS-like distribution (without SMBHB signal), we found that the posterior might be in tension with those constraints. The loop simulation bound also seems to constrain the $(N_{c},N_{k})=(2,0)$ scenario both for the LRS-like and the Model-C-like distributions. Therefore, we may conclude that the AH model could explain the data by itself if the NG-like loops follow LRS model or Model-C with kink-dominated loop spectrum. Otherwise, they should be sub-dominant over the other plausible signals such as the SMBHB signal. In the latter case, SGWB observations becomes not so constraining for the AH model due to the degeneracy of $G\mu$ and $f_{\rm NG}$ in determining the signal amplitude. To further investigate the model, one should combine the constraint from the particle radiation as discussed in Refs. [58, 57, 2], which can place the upper bounds on $G\mu$ depending on the interaction between the AH sector and the other sectors.

### v3

@@ Abstract
In a recent paper, the NANOGrav collaboration studied new physics explanations of the observed pulsar timing residuals consistent with a stochastic gravitational wave background (SGWB) [1], including cosmic strings in the Nambu-Goto (NG) approximation. Analysing one of current models for the loop distribution, it was found that the cosmic string model is disfavored compared to other sources, for example, super massive black hole binaries (SMBHBs). When both SMBHB and cosmic string models are included in the analysis, an upper bound on a string tension $G\mu\lesssim 10^{-10}$ was derived. However, the analysis did not accommodate results from cosmic string simulations in an underlying field theory, which indicate that at most a small fraction of string loops survive long enough to emit GW. Following and extending our previous study [2], we suppose that a fraction $f_{\rm NG}$ of string loops follow NG dynamics and emit only GWs, and study the three different models of the loop distribution discussed in the LIGO-Virgo-KAGRA (LVK) collaboration analyses. We re-analyse the NANOGrav 15yrs data with our signal models by using the NANOGrav ENTERPRISE analysis code via the wrapper PTArcade. We find that loop distributions similar to LVK Model B and C yield higher Bayes factor than Model A analysed in the NANOGrav paper, as they can more easily accommodate a blue-tilted spectrum of the observed amplitude. Furthermore, because of the degeneracy of $G\mu$ and $f_{\rm NG}$ in determining the signal amplitude, our posterior distribution extends to higher values of $G\mu$, and in some cases the uppermost value of credible intervals is close to the Cosmic Microwave Background limit $G\mu\lesssim 10^{-7}$. Hence, in addition to the pulsar timing array data, further information about the fraction of long-lived loops in a cosmic string network is required to constrain the string tension.
@@ 1 Introduction
Recently, PTA collaborations (NANOGrav, EPTA/InPTA, PPTA and CPTA) have reported evidence for an isotropic SWGB as the expected Hellings-Downs (HD) angular correlation between pulsars’ line of sight were measured at $3-4\sigma$ confidence level [20, 21, 22, 23]. With high significance of the common spectrum and the strong evidence of HD correlation, it is highly motivated to investigate whether current observations can be explained by SGWB from the population of cosmic string loops [1, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40] (see also Refs. [41, 42] for global strings.). Indeed, NANOGrav and EPTA performed Bayesian inference for the models of NG loop distribution [43, 44], whose SGWB spectrum is controlled by the dimensionless string tension $G\mu$. Because the NG models have difficulty reproducing the blue-tilted spectrum favored by the data with consistent amplitude, upper bounds on the string tension are derived when combining the NG models and the “more favored” super massive black hole binary (SMBHB) signal [1, 24].
@@ 2.1 SGWB from field theory string loops
As mentioned in the introduction, long-lived oscillating NG-like loops, required for a significant GW signal, have not yet been observed in large-scale field theory simulations [53]. Therefore their length distribution, needed for calculations of the GW power spectrum, is completely unknown. In Ref. [53], (a part of) this uncertainty in their distribution was parameterised by allowing a fraction $f_{\rm NG}$ of loops to survive to radiate only gravitationally. Subsequently in Ref. [2], it was assumed that all the AH loops would have the same length distribution as in an NG network ${\sf n}(l,t)$, and hence that the distribution of NG-like loops would be $f_{\rm NG}{\sf n}(l,t)$. Then the SGWB from NG-like distribution in the AH string network is quantified as
@@
We follow this quantification and discuss possible models of NG loop distributions in the rest of the section.
We emphasise that use of the models of NG loop distribution is only a first approximation since there is no information from the AH simulations what ${\sf n}(l,t)$ should be used. In fitting the NANOGrav data with Eq. (2.1), the interpretation of $f_{\rm NG}$ as parameterising the fraction of NG-like loops holds only if the NG models are close to the true distribution of any loops in a field theory network. This closeness is our underlying assumption in the following discussion. In practice, the normalisation of ${\sf n}(l,t)$ is highly affected by the lengths of the long-lived loops at their production [65] and also by the efficiency of loop production [66]. Therefore, even if the functional form of ${\sf n}(l,t)$ is correct, such an ambiguity in the normalisation might be still encapsulated in $f_{\rm NG}$. Nevertheless, assuming that the loop size at the production well agrees with the BOS model, the authors of Ref. [66] quantify the suppression of the loop production efficiency in the AH string network and the resultant overall amplitude of SGWB due to the classical radiation, which was only by a factor $\sim$ 2. With the (optimistic) expectation that the correction in Eq. (2.1) is O(1), the development of improved models of loop distribution is left to future work. In any case, the difference in the following results for different reference NG loop models should be understood as the underlying uncertainty in the true loop distribution.
@@ 2.2 SGWB from NG string loops
with $H(z)$ and $t(z)$ being the Hubble parameter and cosmic time at redshift $z$. The function $C_{n}(f)$ gives the number density of loops emitting GWs observed at frequency $f$ in the $n$-th harmonic, while the average loop power spectrum $P_{n}$ represents the average GW power emitted by the $n$-th harmonic of a loop. We note that if there is a constant factor uncertainty common to all $P_{n}$, then it would be absorbed by $f_{\rm NG}$ in (2.1) and affect the interpretation of the analyses. For simplicity, we assume that $f_{\rm NG}$ is solely contributed from the uncertainty in the NG-like loop distribution in the following section.
The constants $P_{n}$ depend on the average shape of the loops, and in particular the number of cusps per oscillation and the number of kinks travelling around the loop, which dictate the high frequency behaviour. Their sum (or the so-called decay constant of strings) $\Gamma$ can be decomposed into three contributions as
@@ 2.2.1 BOS model (LVK-Model A)
As a power spectrum of loops $P_{n}$, the “smoothed” model [67] is often adopted for this BOS distribution (see e.g. the analysis of Refs. [68, 20]. It is constructed from the numerical simulation taking into account the gravitational backreaction and different from the simple decomposition in Eq. (2.5). Following the results of their simulation, we set $\Gamma=50$ and extract the values of $P_{n}$ from the Figure 3 and 4 of Ref. [67]. As another example in this class of power spectra, we also consider the kink-dominated case $(N_{c},N_{k})=(1,100)$ with Eq. (2.5), which yields $\Gamma=710$. Notice that the value of $\Gamma$, which affects the amplitude of SGWB, is totally different from those in Ref. [67]. Given the uncertainty on the initial number of kinks at the loop production, however, this case was analysed in Refs. [19, 69, 21]. Therefore, these choices allow for a direct comparison of our results below with those of previous analyses.
@@ 2.2.2 LRS model (LVK-Model B)
In the context of recent PTA observation, this model was analysed (together with the BOS model) by the EPTA collaboration with their second data release [24]. Although they do not perform a model selection analysis, the posterior distribution they obtained somewhat supports the NANOGrav result mentioned above. That is, when they include the SMBHB signal in their analysis, they observe the posterior distribution for $G\mu$ extends downward to the limit of the prior (see the right panel of Fig. 8 in App. A or Fig. 15 of Ref. [24]). This indicates that with the addition of the SMBHB signal, which is more preferred by the data, the magnitude of string signal (or $G\mu$) could only be bounded above.
Following the analyses performed in Refs. [69, 24], here we consider two different scenarios for the loop power spectrum. The first one is $(N_{c},N_{k})=(2,0)$, which yields $\Gamma=57$, a value close to that observed in the simulations. Another one is again $(N_{c},N_{k})=(1,100)$ allowing for the large number of kinks. Here again, our choice is intended as a comparison with the results of those previous studies. If the smoothed model is adopted instead of the first model, qualitatively similar results are expected due to the dominance of cusp and the close value of $\Gamma$ (see Ref. [68] for comparison of different models of loop power spectrum).
@@ 3.1 Summary of the analyses
As discussed in Refs. [1, 24], however, the information carried by the “signal” is not particularly constraining at the moment and the result should crucially depend on the priors. Therefore, especially under the significant uncertainty in the underlying AH loop distribution, too much emphasis should not be placed on the exact number of Bayes factors reported in this study. Nevertheless, it can be considered as a good indicator for comparing which of the models we analysed is relatively better at explaining the data.
@@ 3.2.1 $f_{\rm NG}$-BOS models
We also analyse the case where the indicated SGWB signal is the superposition of the signal from SMBHB and the AH strings. In Fig. 3, the posterior distribution of the string parameters and SMBHB is plotted. Hereafter in the posterior plots, gw-bhb-0 and gw-bhb-1 represent $A_{\rm BHB}$ and $\gamma_{\rm BHB}$ respectively. Again, the left panel is for the smoothed loop power spectrum and the right panel is for the kink dominated spectrum. In both cases, it can be seen that the data favours the sub-dominance of the string signal and the credible intervals of $(G\mu,f_{\rm NG})$ are bounded from above. As expected, the upper most value of $G\mu$ in the credible interval becomes larger than the case with pure BOS model $[-11.99,-9.90]$ (see Table 4 in Ref. [1]). In other words, higher $G\mu$ (or symmetry breaking scales) are to some extent still allowed for by taking into account that not all string loops emit GWs.
@@ 4 Discussion
In quantifying the SGWB signal from long-lived AH string loops, here we took three representative models for NG loop distribution [43, 44, 71]. These models are only approximations since there are no long-lived loops observed in simulations of networks the AH model. Therefore, the interpretation of $f_{\rm NG}$ as a fraction of long-lived loops and the constraints on $(G\mu,f_{\rm NG})$ derived in the analyses physically make sense only if the model distribution is close to the actual loop distribution. In this sense, the adoption of the three models is a case-by-case examination of three distinct possibilities, each assuming that the real distribution of long-lived loops matches that of a particular NG model. Further study of possible NG-like loop production is indispensable for the refinement of SGWB constraint on the AH model: the current status is only that $f_{\rm NG}\lesssim 0.1$ and no information on ${\sf n}(l,t)$.
@@
Our analyses were based on the wrapper PTArcade [61, 62] and conducted in the similar way to the NANOGrav study of new physics [1]. Note that while the BOS model was analysed as a NG loop distribution model and subjected to the model selection analysis by NANOGrav [1], the LRS and LVK-Model-C were not (see Ref. [24] for the posterior for the LRS model derived from EPTA data). As summarized in Sec. 3, we performed Bayesian inference and model selection analysis for all three models assuming only the signal from strings and also assuming the signal from SMBHB. We also investigated the different loop power spectra for each model. In all cases our posterior for the AH model extends to a higher value of $G\mu$ (with a decrease in the value of $f_{\rm NG}$), compared to the pure NG model results in Refs. [1, 24]. Phenomenologically, this suggests that symmetry breaking in the AH model is still possible on relatively high energy scales. It is also worth mentioning that for all the loop distribution models, the posterior of SMBHB parameters become more consistent with the theory-motivated prior used in the NANOGrav paper [20], when the string signal is superimposed to the SMBHB signal. When comparing each model, the LRS-like and model-C-like distribution overall had larger Bayes factors (against the SMBHB signals) than the BOS-like distribution. This is due to the fact that for a wider range of $G\mu$ values, loops of the LRS model and LVK model-C produce a blue-tilted SGWB, which is favored by the data. Note again that the differences in our results for different reference NG loop models should be regarded as an estimate of the uncertainty in the AH loop distribution.
@@
After deriving the posterior distributions, we compared them to the existing bounds such as $G\mu\lesssim 10^{-7}$ from CMB [78, 56] and $f_{\rm NG}\lesssim 0.1$ derived from the AH loop simulation [53]. For the BOS-like distribution (without SMBHB signal), we found that the posterior might be in tension with those constraints. The loop simulation bound also seems to constrain the $(N_{c},N_{k})=(2,0)$ scenario both for the LRS-like and the Model-C-like distributions. Therefore, we may conclude that the AH model could explain the data by itself if the NG-like loops follow LRS model or Model-C with kink-dominated loop spectrum. This may cast a doubt on the AH model as the origin of the SGWB becoming visible at the PTAs, given that the validity of LRS model has been recently questioned in terms of energy transfer (see e.g. Ref. [79]). Otherwise, they should be sub-dominant over the other plausible signals such as the SMBHB signal. In this case, SGWB observations becomes not so constraining for the AH model due to the degeneracy of $G\mu$ and $f_{\rm NG}$ in determining the signal amplitude. To further investigate the model, one should combine the constraint from the particle radiation as discussed in Refs. [59, 58, 2], which can place the upper bounds on $G\mu$ depending on the interaction between the AH sector and the other sectors.